Calculate the following binomial probability by either using one of the binomial probability tables, software, or a calculator using the formula below. Round your answer to 3 decimal places. P(x | n, p) = n! (n − x)! x! · px · qn − x where q = 1 − p P(x = 10, n = 12, p = 0.75) =
Answer #1

Solution :

Given that ,

p = 0.75

q = 1 - p = 1 -0.75 = 0.25  

n = 12

x = 10

Using binomial probability formula ,

P(X = x) = ((n! / (n - x)!) * px * qn - x

P(X =10) = ((12! / (12 - 10)!) * 0.7510* 0.25 12- 10

= 0.2323

Probability = 0.232